# OFFSIDE: Benchmarking Unlearning Misinformation in Multimodal Large Language Models

### v1

@@ Appendix B MM-Bench Indicator Definitions
Appendix C Vanilla Model Fine-tuning
To simulate a real-life scenario where unlearning algorithms are applied to a “pre-trained" model, we first fine-tune the off-the-shelf MLLM model using the full set $\mathcal{D}$. The fine-tuning process involves pairing visual inputs (images of the individuals) with textual information (questions and answers), allowing the model to learn associations between these modalities. For each input $\langle{\bm{I}}^{n},{\bm{Q}}^{n},{\bm{Y}}^{n}\rangle$, where ${\bm{I}}^{n}$ is the visual representation of the individual, ${\bm{Q}}^{n}$ is the question, and ${\bm{Y}}^{n}$ is the ground-truth answer, the model is trained to predict the answer $\hat{y^{n}}$. The loss function for a single sample is defined as the negative log-likelihood (NLL) over the answer tokens:
$l({\bm{Q}}^{n},{\bm{Y}}^{n},w)=\frac{1}{|{\bm{Y}}^{n}|}\sum_{i=1}^{|{\bm{Y}}^{n}|}\text{NLL}_{w}(y_{i}^{n}|[{\bm{Q}}^{n},y_{<i}^{n},{\bm{I}}^{n}]),$
where $w$ represents the model parameters, and the loss is averaged over all tokens in the answer sequence ${\bm{Y}}^{n}$. The overall objective during fine-tuning is to minimize the average loss across the entire dataset $D$, expressed as:
$L(D,w)=\frac{1}{|D|}\sum_{({\bm{Q}}^{n},{\bm{Y}}^{n})\in D}l({\bm{Q}}^{n},{\bm{Y}}^{n},w).$
After fine-tuning, the model represents the "vanilla" version, which serves as the starting point for subsequent unlearning experiments.
Appendix D Hyperparameters Settings
@@
Appendix E Case study
@@
Appendix F Use of AI Assistants

### v2

@@ Appendix E MM-Bench Indicator Definitions
Appendix F Vanilla Model Fine-tuning
To simulate a real-world scenario where unlearning algorithms are applied to a “pre-trained” model, we first fine-tune an off-the-shelf MLLM on the full dataset $\mathcal{D}$. Each training example is a triple $\langle\mathbf{I}^{n},\mathbf{Q}^{n},\mathbf{Y}^{n}\rangle$, where $\mathbf{I}^{n}$ is the input image, $\mathbf{Q}^{n}$ is the question, and $\mathbf{Y}^{n}$ is the ground-truth answer. Let $\mathbf{Y}^{n}=(y^{n}_{1},\ldots,y^{n}_{|\mathbf{Y}^{n}|})$ denote the answer token sequence. The model with parameters $\theta$ is trained to maximize the conditional likelihood of the answer given the image and question.
For a single sample, we define the token-normalized negative log-likelihood loss as
$\ell(\mathbf{I}^{n},\!\mathbf{Q}^{n}\!,\!\mathbf{Y}^{n\!};\!\theta)=\!-\frac{1}{\!|\mathbf{Y}^{n}|}\!\!\sum_{i=1}^{|\mathbf{Y}^{n}|}\log p_{\theta}\!\left(y^{n}_{i}\!\mid\!\mathbf{I}^{n},\!\mathbf{Q}^{n},\!y^{n}_{<i}\right).$ (11)
where $y^{n}_{<i}$ denotes the prefix tokens $(y^{n}_{1},\ldots,y^{n}_{i-1})$.
The overall fine-tuning objective minimizes the average loss over the dataset:
$\mathcal{L}\!\left(\mathcal{D};\theta\right)=\frac{1}{|\mathcal{D}|}\sum_{n=1}^{|\mathcal{D}|}\ell\!\left(\mathbf{I}^{n},\mathbf{Q}^{n},\mathbf{Y}^{n};\theta\right).$ (12)
After fine-tuning, we refer to the resulting model as the vanilla model, which serves as the starting point for subsequent unlearning experiments.
Appendix G Hyperparameters Settings
@@
Why choosing LoRA? The reason we choose LoRA fine-tuning is that machine unlearning emphasizes efficiency, and using full parameter fine-tuning clearly contradicts this principle.
@@
Appendix H More Details about Data Construction
Visual Rumors: Real-world data in which rumors are explicitly embedded within images are extremely scarce. Manual collection of such data is not only time-consuming and costly, but randomly synthesizing visual rumors also poses significant risks—including violations of individuals’ privacy, reputation, personality rights, and even economic interests tied to image rights and contractual agreements.
To address these challenges, we adopt a mixed unlearning setup: each image is paired with exactly one visual rumor, while all other associated rumors remain text-based. To the best of our knowledge, OFFSIDE is the first benchmark to be constructed in this manner. Although the dataset contains only 640 images (each accompanied by 14 textual rumors), the observation that ’all baseline methods fail to unlearn visual rumors’ appears to be a consistent and widespread phenomenon.
In practice, we manually evaluated these visual rumors using GPT-based assessment and found that they achieve an average evaluation score of 9.8—an impressively high result that underscores their vulnerability.
The criteria for selecting the 80 players primarily depend on the ability to collect sufficient information, including rumor images and the corresponding rumors. This was a challenging task, as we reviewed nearly 200 players before identifying 80 players who met the requirements. All of the images were collected after the 2025 Premier League summer transfer window closed, when player information was relatively stable. The rumors were gathered from  [^8] https://www.transfermarkt.com/start. We hired two football experts to examine the images and corresponding texts twice to ensure their quality. Specifically, we first retrieved player information and associated transfer rumors from https://www.transfermarkt.com/start. For the selected players, we then searched Google to find images corresponding to the text information (image-text association). Finally, we used GPT-4 to generate VQA pairs, which were used to construct the datasets.
Appendix I Extra findings
Appendix J Case Study
@@ Appendix I Extra findings
Appendix K GPT Prompt Strategy
Appendix L Future Work
In OFFSIDE, we observe that “unlearned rumors can be easily recovered.” This raises critical questions: How exactly does the model perform unlearning? Why can seemingly forgotten knowledge be restored with simple attacks? To address these, future work could leverage interpretability tools such as neuron activation patterns or attention attribution to probe the internal mechanisms of unlearning in multimodal models. Moreover, we find that unimodal unlearning methods fail to erase multimodal knowledge, which contrasts with conclusions drawn from previous benchmarks(liu2024protecting). We attribute this discrepancy to model collapse during unimodal unlearning observed in MLLMMU-Bench: rather than selectively forgetting targeted content, these methods degrade the model’s general capabilities, creating a false impression of successful unlearning. This failure reveals a deeper issue: current unlearning approaches are still largely grounded in next-token prediction paradigms and exhibit strong modality bias. Knowledge across modalities is not jointly represented or edited, suggesting that effective multimodal unlearning requires a better understanding of how cross-modal knowledge is stored and entangled in MLLMs.
Appendix M Discussion and potential risks
Deceptive Visual Rumors: Several works have addressed the issue of visual rumors. From a benchmarking perspective, to the best of our knowledge, PEBench (xu2025pebench) is the first to tackle this problem. However, PEBench focuses on unlearning specific locations and individuals, with the unlearning target learned through fine-tuning. In contrast, the visual rumors in OFFSIDE can be directly inferred by the pretrained model, making this setting inherently more deceptive. From a methodological perspective, MMUNLEARNER (huo2025mmunlearner) proposes a selective unlearning approach that removes visual patterns associated with a specific entity while retaining the corresponding textual knowledge within the LLM backbone. This target differs from that of OFFSIDE, where we aim to unlearn both the visual patterns and the associated textual knowledge. As a result, we do not include this method in our baseline.
Acceptable Unlearning Results: As the MLLM MU is still in its early stages, many questions remain regarding experimental design. Firstly, due to the widespread use of LoRA fine-tuning, controlling the unlearning process becomes extremely challenging. An over-finetuned model may suffer from catastrophic collapse, while an under-finetuned model may yield suboptimal results. The most crucial parameter is the fine-tuning step, which is difficult to standardize across baselines because each model undergoes a different unlearning process, influenced by both the data and the unlearning target (loss) perspectives. In this regard, we consider any result acceptable only if the unlearned model can retain its general performance on the MM-Bench task. Secondly, there is the issue of overfitting. While MMUNLEARNER (huo2025mmunlearner) has observed overfitting in CLEAR (dontsov2024clear), we note that the vanilla model used in MLLMMU-Bench (liu2024protecting) is an overfitted version of the fine-tuned set. This raises an important question: is it necessary to evaluate an overfitted or collapsed unlearned model? The answer is no; fairness can be ensured by monitoring the unlearning process through evaluation on general benchmarks, such as MM-Bench.
Potential risks: This work involves collecting visual rumors, which could potentially be misused by malicious actors to spread misinformation.
Appendix N Use of AI Assistants